### v1

# Learning Adaptive and Temporally Causal Video Tokenization in a 1D Latent Space

Abstract
We propose AdapTok, an adaptive temporal causal video tokenizer that can flexibly allocate tokens for different frames based on video content. AdapTok is equipped with a block-wise masking strategy that randomly drops tail tokens of each block during training, and a block causal scorer to predict the reconstruction quality of video frames using different numbers of tokens. During inference, an adaptive token allocation strategy based on integer linear programming is further proposed to adjust token usage given predicted scores. Such design allows for sample-wise, content-aware, and temporally dynamic token allocation under a controllable overall budget. Extensive experiments for video reconstruction and generation on UCF-101 and Kinetics-600 demonstrate the effectiveness of our approach. Without additional image data, AdapTok consistently improves reconstruction quality and generation performance under different token budgets, allowing for more scalable and token-efficient generative video modeling.
Figure 1: AdapTok performs adaptive tokenization both temporally and across samples. Left-to-right shows token allocation adapting over time within a video, while top-to-bottom presents sample-wise allocation under different token budgets. Blue bars indicate the tokens counts used per block.
1 Introduction
The vast amount of video data available in the internet and physical world makes efficient video modeling crucial for building visual-centric agents and world models. With the rise of large language models radford2019language ; brown2020language ; touvron2023llama , auto-regressive (AR) generative modeling has become a universal paradigm for various modalities, including images and videos wang2024emu3 ; wu2024janus ; tian2024visual ; li2024synergen ; jin2024videolavit ; hong2022cogvideo .
Consequently, numerous works gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo ; yan2021videogpt ; ge2022tats ; yu2023magvit ; wang2024larp have begun exploring video tokenization to quantize spatiotemporally continuous video frames into discrete token sequences, enabling autoregressive generation through causal transformers. Unlike image data, video frames exhibit temporal causality despite both being continuous signals. To better model such characteristic, some approaches yu2023magvit ; agarwal2025cosmos ; villegas2022phenaki ; wang2024omnitokenizer have incorporated temporal causal regularization, supporting online streaming processing of video frames during encoding and decoding, thereby significantly improving throughput efficiency.
However, most existing works wang2024omnitokenizer ; agarwal2025cosmos ; wang2024larp encode different frames using a fixed number of tokens, ignoring the inherent redundancy in video data. While some approaches duggal2024alit ; koike2020stochastic ; miwa2025onedpiece ; bachmann2025flextok ; shen2025cat have attempted to encode images with variable-length token sequences, few of them is extended to video data. A recent work named ElasticTok yan2024elastictok , to the best of our knowledge, made an initial attempt to encode different frames within the same video using varying numbers of tokens. However, due to its lack of global planning capabilities and constraints from 2D spatial priors, its token allocation strategy still suffers from imbalances across samples and spatial regions, resulting in suboptimal performance.
From our perspective, an efficient video tokenizer should encompass the following three key characteristics: (1) temporal causality, where the encoding and decoding of preceding frames are independent of subsequent frames, thus supporting online streaming processing; (2) 1-D latent token space, where the token allocation is decoupled from spatial structure, ensuring information density is evenly distributed in space; and (3) adaptive token allocation that flexibly adjusts token numbers based on the information of different samples under a given token budget, achieving global optimality.
Motivated by these considerations, we propose AdapTok, a transformer-based framework consisting of a VQ-tokenizer and a causal scorer. During tokenization, the 3D video frame blocks are transformed into 1D latent tokens using a set of learnable latent tokens and a causal block mask is introduced to enable temporal causal modeling. To enable representing videos with variable length latent tokens, we randomly drop tokens at the tail of each block during training. Additionally, a block causal scorer is also trained to model the reconstruction quality of video frames when using token sequences with different token numbers. During inference, we further propose an adaptive token allocation strategy based on integer linear programming (ILP) named IPAL. Given a desired token count, IPAL adjusts token allocation across different samples based on the predicted metrics by trained scorer, thereby optimizing overall reconstruction quality.
Extensive experiments are conducted to validate the effectiveness of our method. Notably, on UCF-101, AdapTok achieves a video reconstruction performance of FVD=28, significantly outperforming existing methods. With our adaptive token allocation strategy, AdapTok achieves Pareto optimality between performance and token count. As shown in Fig.1, AdapTok can adaptively allocate tokens based on video frame content, both temporally and sample-wise, and the reconstruction performance continues to improve as the number of tokens increases. Furthermore, we evaluated video generation performance using AdapTok as a tokenizer on UCF-101 class-conditional generation and Kinetics-600 frame prediction tasks, where our method also demonstrated significant performance improvements. Comprehensive ablation studies are also conducted to verify the necessity of each component.
Our contributions can be summarized as follows:
  • We propose AdapTok, an adaptive framework designed for video tokenization that simultaneously features temporal causality, a 1D latent token space, and a flexible adaptive token allocation strategy, yielding more compact token representations.
  • AdapTok features a novel causal scorer and corresponding adaptive token allocation strategy that dynamically adjusts token allocation across different samples during inference, optimizing overall reconstruction quality.
  • AdapTok significantly improves existing video reconstruction performance, achieves Pareto optimality between token allocation quantity and reconstruction performance, and enhances video generation quality of AR models on class-conditional generation and frame prediction tasks.
2 Related Work
2.1 Discrete Video Tokenizer
Autoregressive (AR) transformers have emerged as powerful generative models across domains, including vision. To handle continuous visual signals, VQ-VAE van2017vqvae and its extensions esser2021vqgan ; yu2021vitvqgan ; zheng2023online ; zhang2023regularized ; mentzer2023fsqvqvae introduce vector quantization to convert images into discrete code sequences. When applied to videos, tokenization methods have evolved significantly. Early approaches use per-frame tokenization gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo , where image tokenizers are directly applied without temporal modeling. Block-based tokenization yan2021videogpt ; ge2022tats ; yu2023magvit improves on this by using 3D convolutions to encode short clips with local spatio-temporal context. More recently, 1-D tokenizers wang2024larp introduce learnable holistic tokens to capture global information across the entire video. Despite these advances, these methods are inherently non-causal, relying on full access to the video during encoding, which limits their applicability to streaming or autoregressive generation tasks.
To address the limitations of non-causal tokenizers, several works have explored causal temporal tokenization architectures. One line of work adopts causal convolutions for autoregressive modeling yu2023magvitv2 . Building on this, Cosmos-Tokenizer agarwal2025cosmos introduces both causal convolution layers and causal temporal attention layers for temporal modeling. In parallel, another line of work focuses on transformer-based architectures villegas2022phenaki ; wang2024omnitokenizer , where causality is enforced via attention masks, enabling autoregressive modeling using only past frames. Following this trend, our method also adopts a transformer-based causal architecture, facilitating better adaptation to downstream Physical AI applications that require a temporal causal setting.
2.2 Adaptive Tokenizer
Recent works have explored adaptive tokenization to allow variable length representations. For image tokenization, ALIT duggal2024alit dynamically adjusts the token count by recursively distilling 2D image tokens into 1D latent sequence. Another common strategy is “tail drop” koike2020stochastic , which applies higher dropout rate on the tail of the latent feature to learn ordered representations. Building on this idea, One-D-Piece miwa2025onedpiece and FlexTok bachmann2025flextok introduce tail-drop regularization for 1D image tokenizers. To support adaptive inference, CAT shen2025cat predicts image complexity using captions and large language models (LLMs), assigning each input to one of three compression ratios (8$\times$, 16$\times$, 32$\times$).
Extending adaptive tokenization to videos, ElasticTok yan2024elastictok applies tail dropping at the end of each block and supports adaptive inference by selecting token counts via a fixed threshold. However, due to the local spatial-temporal nature of these tokens—rather than forming a unified 1D sequence—it is difficult to gather informative content at the head, limiting the effectiveness of tail dropping. Moreover, relying on a fixed threshold prevents token allocation under a specified global budget. In contrast, our AdapTok adopts a 1D transformer that decouples token allocation from spatial structure. We further introduce an online scorer with IPAL for efficient token allocation, to enable globally optimal token allocation under a given token budget.
3 Method
This work aims to present a flexible video tokenizer that adaptively tokenizes videos both temporally and sample-wise. In pursuit of this goal, we highlight three core designs: an adaptive video tokenizer which allows for representing videos in variable-length (Sec. 3.1), a scoring module (Sec. 3.2), and a token allocation strategy (Sec. 3.3). Complementing this, we also explore its integration in video generation (Sec. 3.4). The overall framework of our method is illustrated in Fig. 2.
Figure 2: Overview of the proposed AdapTok framework. (a) Adaptive Tokenizer: composed of a block-causal encoder, a block-wise mask sampler, and a block causal decoder for reconstructing video from adaptively masked latent representations. (b) Attention Masks: block-causal attention patterns used in the encoder, decoder, and scorer. (c) Adaptive Scorer: top-down illustrates the generation of ground-truth scores - duplicated latent tokens $z$ are masked and decoded into videos, and perceptual loss $\mathcal{L}_{P}$ is computed as the quality scores $s$. Bottom-up shows the scorer predicting block-wise quality scores $\hat{s}$ from continuous latents $z$ and quantized latents $z_{q}$.
3.1 Adaptive Tokenizer
3.1.1 3D Patchification
Given a video input $x\in\mathbb{R}^{T\times H\times W\times 3}$, where $T$ is the number of frames and $H\times W$ is the spatial resolution, we split it into non-overlapping spatio-temporal patches with a patch size of $t\times p\times p$. The video patches are then flattened and projected into patch embeddings $e\in\mathbb{R}^{L\times d}$, where $L=\frac{T}{t}\times\frac{W}{p}\times\frac{H}{p}$ denotes the total number of tokens. The $L$ tokens are then divided into $K$ blocks.
3.1.2 Block Causal Transformer
Block Causal Encoder. To design an adaptive tokenizer, it is crucial to avoid tokens that explicitly correspond to local spatial-temporal regions. To this end, the encoder employs a set of learnable latent tokens to transform the patch embeddings into a 1-D sequence. Specifically, the latent tokens $q_{\mathrm{enc}}\in\mathbb{R}^{N\times d}$ are concatenated with image tokens $e\in\mathbb{R}^{L\times d}$, forming a combined input of $(L+N)$ tokens to a block-causal transformer encoder $\mathcal{E}$. The output corresponding to the latent tokens, denoted as $\tilde{z}=\mathcal{E}(e\oplus q_{\mathrm{enc}})_{L:(L+N)}$, serves as the learned latent representation.
Across blocks, block-causal attention is applied: all tokens in a block (both latent and image tokens) can only attend to tokens from the same or previous blocks, thereby enforcing a causal dependency structure. The pattern for the encoder attention mask is shown in Fig. 2(b-i).
Block-mask Sampler. We train AdapTok to learn variable length codes over a short video block (e.g., 4 frames) by randomly dropping the tail codes for each block during training. Specifically, for each latent block $\tilde{z}_{i}$, we randomly sample the number of tokens to retain, denoted as $\ell_{i}$. After sampling $\{\ell_{i}\}_{i=1}^{K}$, a binary latent mask $m\in\{0,1\}^{M\cdot K}$ is constructed, where the first $\ell_{i}$ values in each of the $K$ blocks are set to 1, and $M$ is the total length of each block that satisfies $N=M\cdot K$.
$m^{\prime}=\ [m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\ ],\ m_{i}=\ [\mathbbm{1}_{j\leq\ell_{i}}]_{j=1}^{M}.$ (1)
The mask $m^{\prime}$ is then applied to the latent token sequence $\tilde{z}$, where tokens with $m_{j}^{\prime}=0$ are dropped during training. The resulting sequence, denoted as $z=\mathrm{BlockTailDrop}(\tilde{z}\ |\ m^{\prime})$, is used for decoding, enabling the model to learn efficient, variable-length representations.
Quantizer. Following wang2024larp , a stochastic vector quantization (SVQ) is adopted to the quantizer $\mathcal{Q}$. Similar to VQ, a codebook $C\in\mathbb{R}^{c\times d^{\prime}}$ containing $c$ codes is maintained. Given a visual feature $z$, the cosine similarity with all code vectors in $C$ is computed, followed by a softmax to obtain a probability distribution. An index $x_{\mathrm{Ind}}$ is then sampled from this distribution via a categorical distribution:
$x_{\mathrm{Ind}}\sim\mathrm{Categorical}\left(\mathrm{softmax}\left(\left\{\frac{v\cdot C_{i}}{\|v\|\|C_{i}\|}\right\}_{i=1}^{c}\right)\right).$ (2)
Given the sampled index $x_{\mathrm{Ind}}$, the quantized feature $z_{q}$ is retrieved via codebook lookup, i.e., $z_{q}=C_{x_{\mathrm{Ind}}}$. To enable differentiable training, the straight-through estimator bengio2013estimating is applied.
Block Causal Decoder. The quantized tokens are then concatenated with a set of learnable image latent tokens $q_{\mathrm{dec}}\in\mathbb{R}^{L\times d}$, and the combined $(N^{\prime}+L)$ tokens are fed into a block-causal decoder $\mathcal{D}$, implemented as a transformer. The last $L$ outputs, denoted as $\hat{z}=\mathcal{D}(q_{\mathrm{dec}}\oplus z_{q},\mathcal{M}^{\prime})_{N^{\prime}:(N^{\prime}+L)}$, are used to reconstruct the video $\hat{x}\in\mathbb{R}^{T\times H\times W\times 3}$. Here, $N^{\prime}$ denotes the number of latent tokens remaining after tail-drop, and $\mathcal{M}^{\prime}$ is the attention mask applied in the decoder.This mask is derived by applying the dropout mask $m^{\prime}$ to the original block-causal attention pattern $\mathcal{M}$, as shown in Fig 2(b-ii),
$\mathcal{M}^{\prime}=\mathrm{BlockTailDrop}(\mathcal{M}\ |\ m^{\prime}).$ (3)
Within this mask, quantized latent tokens $z_{q}$ attend only to quantized tokens from the same or preceding blocks, while image latent tokens $q_{\mathrm{dec}}$ attend to all image tokens within the same block and to quantized tokens from the same or earlier blocks.
3.1.3 Training
Following previous works esser2021vqgan ; ge2022tats ; wang2024larp , the tokenizer is optimized with a composite objective:
$\mathcal{L}=\mathcal{L}_{R}+\mathcal{L}_{VQ}+\mathcal{L}_{P}+\mathcal{L}_{G}+\mathcal{L}_{prior},$ (4)
where $\mathcal{L}_{R}$ denotes the $L_{1}$ reconstruction loss, $\mathcal{L}_{VQ}$ the quantizer loss, $\mathcal{L}_{P}$ the perceptual loss zhang2018unreasonable , $\mathcal{L}_{G}$ the adversarial loss goodfellow2014generative to enhance visual quality, and $\mathcal{L}_{prior}$ the autoregressive prior loss wang2024larp for modeling the latent sequence.
3.2 Adaptive Scorer
To enable adaptive token allocation during inference, we introduce a block-causal scorer that learns to predict reconstruction quality under different token budgets, as illustrated in Fig. 2(c).
Generating Ground-Truth Scores. At each iteration, a target block index $q\sim\mathcal{U}(0,K-1)$ is sampled. We duplicate the latent sequence ${z}$ and apply a series of latent masks $\{m^{\prime}_{p}\}$, each corresponding to a different number of tokens $\ell_{q}$, ranging from the supported range of encoding lengths. The token lengths $\ell_{i<q}$ for preceding blocks are randomly sampled, while all subsequent blocks are fully masked, i.e., $\ell_{i>q}=0$. For the $p$-th duplicated sample, the latent mask is given by:
$m_{p}^{\prime}=\ [m_{p,1}\oplus m_{p,2}\oplus\cdots\oplus m_{p,q}\oplus\mathbf{0}\ ],\ m_{p,i<q}=\ [\mathbbm{1}_{j\leq\ell_{i}}]_{j=1}^{M},\ m_{p,q}=\ [\mathbbm{1}_{j\leq p}]_{j=1}^{M}.$ (5)
For each $m^{\prime}_{p}$, the perceptual loss $\mathcal{L}_{P}$ over the $q$-th block is computed, serving as the quality scores $s$.
Block Causal Scorer. The scorer $\mathcal{S_{\phi}}$, implemented as a transformer encoder with block causal attention, takes both the continuous latent tokens $z$ and the quantized tokens $z_{q}$ as input. In a single forward pass within a block, it predicts quality scores for all candidate token lengths $\ell_{q}$, denoted as:
$\hat{s}=\mathcal{S_{\phi}}(z\oplus z_{q},\mathcal{M}^{\prime}_{s})_{qM:(q+1)M},$ (6)
where $\mathcal{M}^{\prime}_{s}$ is a block-causal attention mask derived from a base pattern $\mathcal{M}_{s}$ (Fig 2(b-iii)) by applying a dropout mask $m^{\prime}_{s}$. The dropout mask $m^{\prime}_{s}$ is constructed in a manner similar to $m^{\prime}_{p}$ (used for ground-truth generation), except that $\ell_{q}$ is fixed to the maximum block token length $M$.
The mean squared error (MSE) loss $\mathcal{L}_{s}$ is computed between the predicted score $\hat{s}$ and the ground-truth quality scores $s$.
3.3 Adaptive Token allocation
At inference, we introduce an integer programming-based adaptive allocation strategy (IPAL) that assigns a variable number of tokens to each video block, aiming to optimize the overall reconstruction quality under a global token budget, as detailed in Algorithm 1. Specifically, for each block in a mini-batch $B$, the block-causal scorer $\mathcal{S}_{\phi}$ predicts a vector of quality scores $\hat{s}_{k}$ for each sample $k$, across candidate token lengths. A binary variable $b_{kj}\in\{0,1\}$ indicates whether sample $k$ is assigned $j$ tokens. The ILP is formulated as:
$\min_{b}\sum_{k,j}\hat{s}_{kj}\cdot b_{kj}\quad\text{s.t.}\quad\sum_{j}b_{kj}=1\ \forall k,\quad\sum_{k,j}j\cdot b_{kj}=B\cdot N_{b},$ (7)
where $B$ is the batch size and $N_{b}$ is the average number of tokens to be assigned per block.
The objective minimizes the total predicted scores $\hat{s}_{kj}$ (perceptual loss) across a mini-batch, weighted by the binary assignment variables $b_{kj}$. The first constraint ensures that each sample selects exactly one token length, while the second constraint enforces the total token budget across the batch. The optimal assignment $b^{*}$ yields the final token count for each sample: $n_{k}=\sum_{j}j\cdot b^{*}_{kj}$.
1 Inputs: Video $x$, Block idx $q$;
2 Hyperparameters: Average Token $N_{b}$, tokens per block $M$, batch size $B$;
3 $z=\mathcal{E}(x)$, $z_{q}=\mathcal{Q}(z)$;
4 $\hat{s}=\mathcal{S}_{\phi}(z,z_{q})_{qM:(q+1)M}$;
// solve the ILP
5 Define: $b_{kj}\in\{0,1\}$ for binary score bins;
6 Objective: $\mathcal{L}=\min_{b}\sum_{k,j}\hat{s}_{kj}b_{kj}$;
7      s.t. $\sum_{j}b_{kj}=1,~{}\forall k$;
8         $\sum_{k,j}j\cdot b_{kj}=B\cdot N_{b}$;
9 $b^{*}=\text{ILPSolve}(\mathcal{L})$;
10 $n_{k}=\sum_{j}j\cdot b^{*}_{kj}$;
11 Return: assigned tokens $n_{k}$;
Algorithm 1  Integer Linear Programming
Figure 3: Comparison of rFVD by Token Length. AdapTok achieves better rFVD with fewer tokens than existing causal tokenizers.
3.4 Video Generation
Given a sequence of adaptively allocated tokens from the scorer with IPAL, an end-of-block token <EOB> is appended to the end of each block. The resulting concatenated sequence is then fed into a Llama-style transformer touvron2023llama ; touvron2023llama2openfoundation for autoregressive generation. Formally, let $y=(y_{1},y_{2},\cdots,y_{S})$ denote the multi-block adaptive sequence. The model is trained to model the probability distribution by maximize the likelihood of each token $y_{i}$ autoregressively using a standard cross-entropy loss:
$\mathcal{L}=-\sum^{S}_{i=1}\log P(\hat{y}_{i}|c,y_{1:i-1};\theta),$ (8)
where $c$ denotes the conditions, e.g., class labels for class-conditional generation or context tokens for frame prediction, and $\theta$ represents the trainable parameters of the transformer.
4 Experiments
4.1 Experimental Setups
Datasets. We train the model and evaluate its performance on two standard datasets: video reconstruction and class-conditional generation on UCF-101 soomro2012ucf101 and frame prediction on Kinetics-600 carreira2018k600 with 5-frame conditioning.
Implementation details. Following previous works ge2022tats ; yu2023magvitv2 , we adopt $16\times 128\times 128$ video clips for both training and evaluation across all tasks. Patch embeddings are extracted with a patch size of $4\times 8\times 8$, resulting in $L=1024$ tokens per video. A sequence of latent tokens $q_{\mathrm{enc}}$ is used to extract these patch embeddings, and is partitioned into $K=4$ blocks, each containing $M=512$ tokens. These latent tokens are then quantized using a codebook $C\in\mathbb{R}^{8192\times 16}$.
For adaptive token sampling, we employ a block-wise mask sampler. In each block, the number of tokens is sampled from a truncated Gaussian distribution with mean $\mu=256$, standard deviation $\sigma=128$, and bounded between $M_{\mathrm{min}}=32$ and $M_{\mathrm{max}}=512$.
AdapTok is trained on UCF-101 and Kinetics-600 for 250 epochs with a batch-size of 128. Adam kingma2014adam is adopt as the optimizer with hyperparameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$. The learning rate is linearly warmed up to $10^{-4}$ and subsequently decayed to $10^{-6}$ using a cosine scheduler.
Fréchet Video Distance (FVD) thomas2018fvd is adopt as the main evaluation metric for both reconstruction and generation. Additionally, we report MSE, PSNR and LPIPS for video reconstruction.
4.2 Main Results
Video Reconstruction. We first evaluate the video reconstruction performance of AdapTok on UCF-101, with results summarized in Tab. 2. We can observe that, even with fewer training data and latent tokens, AdapTok significantly outperforms existing causal video tokenizers, achieving an rFVD of 28 with 2048 latent tokens and 36 with 1024 tokens.
To further evaluate efficiency under varying token budgets, we compare AdapTok with several causal tokenizer baselines. In particular, we implement a strong baseline named CausalTok that shares the same block-causal architecture as AdapTok, but applies a standard non-adaptive pipeline. As shown in Fig. 3, AdapTok consistently achieves lower rFVD across different token counts. Remarkably, it achieves an FVD of 60 using only 512 tokens, outperforming most baselines, and achieves comparable performance to CausalTok while using approximately 1.8$\times$ fewer tokens.
Video Generation. For video generation, We use AdapTok to tokenize videos and train a Llama-style model to autoregressively generate token sequences. As reported in Tab. 2, our model achieves competitive performance on both class-conditional generation and frame prediction. Notably, with only 633M parameters, AdapTok-AR achieves a gFVD of 11 on Kinetics-600 and 67 on UCF-101.
Table 1: Comparison of Video Reconstruction FVD on UCF-101. All models use a causal visual tokenizer. Bold and underline indicate the best and the second best performance, respectively.
Method Data size Codebook Tokens rFVD $\downarrow$
OmniTokenizer wang2024omnitokenizer 1.4M 8,192 1,280 42
ElasticTok yan2024elastictok 356M 64,000 1,024 390
ElasticTok yan2024elastictok 356M 64,000 2,048 93
Cosmos-Tokenizer-DV agarwal2025cosmos 100M 64,000 1,280 140
AdapTok (Ours) <0.5M 8,192 512 60
AdapTok (Ours) <0.5M 8,192 1,024 36
AdapTok (Ours) <0.5M 8,192 2,048 28
Table 2: Class-conditional generation results on UCF-101 and frame prediciton results on Kinetics-600. $\ddagger$ denotes using pretrained image tokenizer.
gFVD $\downarrow$
Method Data size Params K600 UCF
Phenaki villegas2022phenaki 465M 1.8B 36.4 /
MAGVIT-AR yu2023magvit 12M 306M / 265
MAGVIT-v2-AR yu2023magvitv2 1.8M 840M / 109
OmniTokenizer wang2024omnitokenizer 1.4M 650M 32.9 191
CogVideo hong2022cogvideo  $\ddagger$ 5.4M 9.4B 109.2 626
Video-LaVIT jin2024videolavit  $\ddagger$ 10M 7B / 281
TATS ge2022tats <0.5M 321M / 332
AdapTok-AR (Ours) <0.5M 633M 11 67
4.3 Ablation Studies
Adaptive training and inference paradigms. AdapTok leverages two adaptive mechanisms: a block-mask sampler for training and an adaptive scorer for inference. The block-mask sampler enables the model to learn from variable token lengths by randomly sampling token counts per block, while the adaptive scorer allocates tokens based on reconstruction quality scores during inference. As shown in Tab. 3a, leveraging both mechanisms results in the best performance, achieving the lowest rFVD of 36.36. The performance gain becomes more pronounced with lower token budgets in inference, highlighting the importance of our adaptive mechanisms for efficient video representation.
Table 3: Ablation studies. (a) Ablation on adaptive training and inference paradigms. (b) Evaluation on different token allocation strategies. (c) Comparison on token allocation strategies.
Tokens Mask Sampler Scorer rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
1024 37.13 25.92 0.111
38.79 25.29 0.122
36.36 25.72 0.114
512 509.95 14.38 0.368
121.88 22.89 0.170
59.96 24.06 0.144
(a) Adaptive paradigms.
Method rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
Fixed 38.79 25.29 0.122
BiThr 42.12 25.23 0.120
BiDelta 38.13 25.65 0.115
ILP 36.36 25.72 0.114
(b) Token allocation strategies.
Metric rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
SSIM 37.12 25.74 0.115
PSNR 36.56 25.84 0.115
MSE 36.97 25.75 0.115
Perceptual 36.28 25.72 0.113
(c) Scoring metrics.
Token allocation strategies. We explore several token allocation strategies: 1) Fixed token allocation, which assigns a fixed token count to all blocks; 2) Score-threshold binary search (BiThr), which binary searches for a global score threshold and selects the minimal token count per block with predicted score below it; 3) Delta-score binary search (BiDelta), which selects the minimal token count where the score improvement between consecutive lengths drops below a binary searched delta; and 4) Integer Linear Programming (ILP), which jointly optimizes token allocation across a mini-batch under an average token budget. As reported in Tab. 3b, the ILP strategy outperforms the other methods across all metrics. For a comparison under varying token lengths, refer to Fig. 4.
Figure 4: Comparison of token allocation strategies. ILP achieves the best performance across four metrics: (a) FVD, (b) PSNR, (c) MSE, and (d) LPIPS.
Token allocation scoring metrics. We also investigate various scoring metrics for the token allocation, including per-sample SSIM, PSNR, MSE, and perceptual loss. As shown in Tab. 3c and Fig. 5, each scoring metric performs best on its corresponding evaluation metric, e.g., PSNR score yields the highest PSNR (Fig. 5b), MSE score leads to the lowest MSE (Fig. 5c). Notably, using perceptual loss as the scoring metric also leads to superior performance on other evaluation metrics, such as rFVD (Fig. 5a), indicating its stronger correlation with overall perceptual quality.
Figure 5: Comparison of scoring metrics. Perceptual loss achieves better overall performance, especially on FVD and LPIPS.
4.4 Visualizations
1D latent token space matters. In Fig. 6, we visualize reconstruction results under different token budgets. Compared to ElasticTok, which relies on local 2D spatial tokens and tends to produce block-like artifacts, AdapTok leverages 1D latent space that enables a coarse-to-fine reconstruction process: early tokens capture global structure, while later tokens refine local details.
Figure 6: Reconstruction comparison with ElasticTok yan2024elastictok under varying token budgets. AdapTok reconstructs videos in a coarse-to-fine manner, where early tokens capture global structure and later ones refine local details, while ElasticTok suffers from blocky artifacts.
Head tokens encode global information. The tail token drop strategy encourages the head tokens to capture more global information. As shown in Fig.7, attention maps reveal that early tokens attend broadly across the entire spatial field, indicating their role in modeling global semantics. In contrast, mid-to-late tokens focus on localized regions, refining finer details.
Figure 7: Attention maps for latent tokens. Each map shows the attention distribution of a token over spatial patches, with tokens ordered from top-left to bottom-right. Early tokens capture global context, while later tokens focus on local details.
5 Conclusion
In this paper, we propose AdapTok, an adaptive video tokenizer designed with temporal causality within a unified 1D latent token space. AdapTok incorporates a novel causal scorer and an ILP-based allocation strategy, IPAL, which dynamically adjusts token usage both temporally and sample-wise during inference. Extensive experiments demonstrate that AdapTok not only achieves superior reconstruction quality compared to existing causal video tokenizers, but also provides a favorable trade-off between performance and token budgets, and improves video generation in both class-conditional and frame prediction tasks.
Limitations and Future work. In this work, we focus on the design of a sample-wise, content-aware, and temporally adaptive tokenization framework. Our current implementation adopts a discrete tokenizer based on VQ-VAEs van2017vqvae ; esser2021vqgan , and future work may explore continuous alternatives to validate the generality of the proposed framework. In addition, our model is trained on less than 0.5M publicly available videos due to limited computational resources. Future efforts will focus on scaling up model capacity and dataset diversity to improve generalization and applicability across broader domains.
Appendix A Implementation Details
The detailed training hyper-parameter settings for the AdapTok, Scorer, AdapTok-AR (class-conditional generation on UCF-101) and AdapTok-FP (frame prediciton on Kinetics-600) are reported in Tab. 4.
Table 4: Hyper-parameters for AdapTok models.
AdapTok Scorer AdapTok-AR AdapTok-FP
Model parameters
[…]
Frame Resolution $16\times 128\times 128$ $16\times 128\times 128$ $16\times 128\times 128$ $16\times 128\times 128$
Patch Size $4\times 8\times 8$ $4\times 8\times 8$ $4\times 8\times 8$ $4\times 8\times 8$
Hidden Size 768 768 1280 1280
Transformer Layers 12 12 30 30
Training
[…]
Learning rate $1e^{-4}$ $1e^{-4}$ $6e^{-4}$ $6e^{-4}$
Beta1 0.5 0.5 0.9 0.9
Beta2 0.9 0.9 0.95 0.95
Weight decay 0 0 0.05 0.05
Scheduler type cosine cosine cosine cosine
Warmup epochs 8 8 4 1
Batch size 128 128 64 64
Epochs 250 20 3000 75
GPUs 32 32 8 16
Training Time 112h 9h 59h 55h
Appendix B Adaptive Inference
In addition to the Integer Linear Programming (ILP) method detailed in the main paper, we provide further details for the other three token allocation methods, including Fixed, BiThr and BiDelta.
Fixed token allocation
uses the same number of tokens across all video samples and blocks. The latent mask is given by:
$m^{\prime}=\ [m_{1}\oplus m_{2}\oplus\cdots\oplus m_{K}\ ],\ m_{i}=\ [\mathbbm{1}_{j\leq N_{b}}]_{j=1}^{M}.$ (9)
where $N_{b}$ is the fixed number of tokens allocated per block.
Score-threshold binary search (BiThr)
binary searches a global score threshold which assigns the minimum token counts that satisfy a desired video quality. Specifically, given a score threshold $s_{i}$, the token counts are assigned by selecting the first position where the score drops below $s_{i}$. The threshold is iteratively updated via binary search until the average token count matches the target value, as detailed in Algorithm 2.
Delta-score binary search (BiDelta)
follows the same binary search procedure as BiThr, but operates on delta scores instead of raw scores. These delta scores, computed using a difference function (see Algorithm 2), quantify the marginal gain in perceptual quality from adding each token. The search aims to find a threshold over these deltas such that the resulting token allocation meets the desired token count.
1 Inputs: Video $x$, Block idx $i$;
2 Hyperparameters: Average Token $N_{b}$, tokens per block $M$, max iterations $K$;
3 $z=\mathcal{E}(x)$, $z_{q}=\mathcal{Q}(z)$;
4 $s=\begin{cases}\mathcal{S}_{\phi}(z,z_{q})_{iM:(i+1)M},&\text{(BiThr)}\\ \mathcal{S}_{\phi}(z,z_{q})_{iM:(iM+M-1)}-\mathcal{S}_{\phi}(z,z_{q})_{(iM+1):(iM+M)},&\text{(BiDelta)}\end{cases}$;
5 $s_{max},s_{min}=\max(s),\min(s)$;
6 for $k=1,\cdots,K$ do
7       $s_{mid}=(s_{max}+s_{min})/2$;
8       $n_{k}=\text{argmax}(s<s_{min})$;
9
10      if $\mathrm{mean}(n_{k})>N_{b}$ then
11            $s_{min}=s_{mid}$;
12      else
13            $s_{max}=s_{mid}$;
14Return: assigned tokens $n_{k}$;
Algorithm 2  Binary Search
Appendix C Visualizations
C.1 Video Reconstruction
Content-aware allocation.
As shown in Fig. 8, AdapTok adaptively allocates tokens based on the visual complexity of each scene. Static or low-motion segments receive fewer tokens, while dynamic or visually complex regions are assigned more, enabling efficient token usage without compromising important content.
Figure 8: AdapTok performs adaptive content-aware tokenization. From the top to the bottom, AdapTok allocates more tokens as scene complexity and motion increase. Blue bars represent the token counts used per block.
Temporal dynamics.
Fig. 9 illustrates how AdapTok adaptively allocates tokens over time. When scene changes, more tokens are assigned to capture the transition, while fewer tokens are used in stable or redundant segments.
Figure 9: AdapTok adaptively allocates tokens based on temporal dynamics. Red boxes highlight the scene transitions, where AdapTok allocates more tokens to capture important temporal changes.
C.2 Video Generation
We present class conditional generation results on UCF-101 in Fig. 10 and frame prediction results on Kinetics-600 in Fig. 11.
Figure 10: Class conditional video generation on UCF-101.
Figure 11: Frame prediction results on Kinetics-600. Red boxes represent the condition frames.
C.3 Token Contribution Analysis
Fig. 12 provides additional attention maps of latent tokens from various samples. These further demonstrate that early tokens tend to attend broadly to capture global context, while mid-to-late tokens focus more on local details.
Figure 12: Additional attention maps for latent tokens. Each map shows how a token attends to spatial patches, ordered from top-left to bottom-right. Early tokens attend broadly to capture global context, while later ones focus on local details.
C.4 Scorer Predictions
Fig. 13 illustrates the scorer’s prediction quality and the effect of increasing token counts on reconstruction. As shown in Fig. 13a, the predicted scores closely match the ground truth scores (perceptual loss) across all blocks. As more tokens are allocated within each block (moving along the x-axis), the score decreases, indicating improved reconstruction quality. This trend is visually confirmed in Fig. 13b, where reconstructions become progressively more accurate as token counts increase.
Figure 13: The visualization of scorer predictions and corresponding reconstruction results. (a) Scorer predictions. Red curves represent the GT scores, while blue denotes the predictions. A total of 2048 tokens are divided into 4 blocks. The x-axis represents the index of the last selected token for each block. (b) Corresponding reconstruction results of frames from the first block under different token counts.

### v2

@@
 Figure 1: AdapTok performs adaptive tokenization both temporally and across samples.
-Left-to-right shows token allocation adapting over time within a video, while top-to-bottom presents sample-wise allocation under different token budgets.
-Blue bars indicate the tokens counts used per block.
+Left-to-right shows token allocation adapting over time, top-to-bottom presents sample-wise allocation under different token budgets.
+Blue bars indicate the tokens counts used per block (i.e., 4 frames).
 1 Introduction
 The vast amount of video data available in the internet and physical world makes efficient video modeling crucial for building visual-centric agents and world models.
-With the rise of large language models radford2019language ; brown2020language ; touvron2023llama , auto-regressive (AR) generative modeling has become a universal paradigm for various modalities, including images and videos wang2024emu3 ; wu2024janus ; tian2024visual ; li2024synergen ; jin2024videolavit ; hong2022cogvideo .
-Consequently, numerous works gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo ; yan2021videogpt ; ge2022tats ; yu2023magvit ; wang2024larp have begun exploring video tokenization to quantize spatiotemporally continuous video frames into discrete token sequences, enabling autoregressive generation through causal transformers.
+With the rise of large language models (Radford et al., 2019; Brown et al., 2020; Touvron et al., 2023a), auto-regressive (AR) generative modeling has become a universal paradigm for various modalities, including images and videos (Wang et al., 2024c; Wu et al., 2024; Tian et al., 2024; Li et al., 2024; Jin et al., 2024; Hong et al., 2023).
+Consequently, numerous works (Gupta et al., 2022; Wu et al., 2022; Hong et al., 2023; Yan et al., 2021; Ge et al., 2022; Yu et al., 2023; Wang et al., 2024a) have begun exploring video tokenization to quantize spatiotemporally continuous video frames into discrete token sequences, enabling autoregressive generation through causal transformers.
 Unlike image data, video frames exhibit temporal causality despite both being continuous signals.
-To better model such characteristic, some approaches yu2023magvit ; agarwal2025cosmos ; villegas2022phenaki ; wang2024omnitokenizer have incorporated temporal causal regularization, supporting online streaming processing of video frames during encoding and decoding, thereby significantly improving throughput efficiency.
-However, most existing works wang2024omnitokenizer ; agarwal2025cosmos ; wang2024larp encode different frames using a fixed number of tokens, ignoring the inherent redundancy in video data.
-While some approaches duggal2024alit ; koike2020stochastic ; miwa2025onedpiece ; bachmann2025flextok ; shen2025cat have attempted to encode images with variable-length token sequences, few of them is extended to video data.
-A recent work named ElasticTok yan2024elastictok , to the best of our knowledge, made an initial attempt to encode different frames within the same video using varying numbers of tokens.
-However, due to its lack of global planning capabilities and constraints from 2D spatial priors, its token allocation strategy still suffers from imbalances across samples and spatial regions, resulting in suboptimal performance.
-From our perspective, an efficient video tokenizer should encompass the following three key characteristics: (1) temporal causality, where the encoding and decoding of preceding frames are independent of subsequent frames, thus supporting online streaming processing; (2) 1-D latent token space, where the token allocation is decoupled from spatial structure, ensuring information density is evenly distributed in space; and (3) adaptive token allocation that flexibly adjusts token numbers based on the information of different samples under a given token budget, achieving global optimality.
+To better model such characteristic, some approaches (Yu et al., 2023; NVIDIA, 2025; Villegas et al., 2022; Wang et al., 2024b) have incorporated temporal causal regularization, supporting online streaming processing of video frames during encoding and decoding, thereby significantly improving throughput efficiency.
+However, most existing works (Wang et al., 2024b; NVIDIA, 2025; Wang et al., 2024a) encode different frames using a fixed number of tokens, ignoring the inherent redundancy in video data.
+While some approaches (Duggal et al., 2024; Koike-Akino & Wang, 2020; Miwa et al., 2025; Bachmann et al., 2025; Shen et al., 2025) have attempted to encode images with variable-length token sequences, few of them is extended to video data.
+A recent work named ElasticTok (Yan et al., 2024), to the best of our knowledge, made an initial attempt to encode different frames within the same video using varying numbers of tokens.
+Nevertheless, due to its lack of global planning capabilities and constraints from 2D spatial priors, its token allocation strategy still suffers from imbalances across samples and spatial regions, resulting in suboptimal performance.
+From our perspective, an efficient video tokenizer should encompass the following three key characteristics: (1) temporal causality, where the encoding and decoding of preceding frames are independent of subsequent frames, thus supporting online streaming processing; (2) 1-D latent token space, where the token allocation is decoupled from spatial structure, ensuring information density is evenly distributed in space; and (3) adaptive token allocation, that adjusts token numbers based on the information of different samples under a given token budget, achieving global optimality.
 Motivated by these considerations, we propose AdapTok, a transformer-based framework consisting of a VQ-tokenizer and a causal scorer.
@@
 As shown in Fig.1, AdapTok can adaptively allocate tokens based on video frame content, both temporally and sample-wise, and the reconstruction performance continues to improve as the number of tokens increases.
-Furthermore, we evaluated video generation performance using AdapTok as a tokenizer on UCF-101 class-conditional generation and Kinetics-600 frame prediction tasks, where our method also demonstrated significant performance improvements.
+Furthermore, we evaluated video generation performance using AdapTok as a tokenizer on UCF-101 class-conditional generation and Kinetics-600 frame prediction tasks, where our method also demonstrated significant performance gains.
 Comprehensive ablation studies are also conducted to verify the necessity of each component.
@@
 Autoregressive (AR) transformers have emerged as powerful generative models across domains, including vision.
-To handle continuous visual signals, VQ-VAE van2017vqvae and its extensions esser2021vqgan ; yu2021vitvqgan ; zheng2023online ; zhang2023regularized ; mentzer2023fsqvqvae introduce vector quantization to convert images into discrete code sequences.
-When applied to videos, tokenization methods have evolved significantly.
-Early approaches use per-frame tokenization gupta2022maskvit ; wu2022nuwa ; hong2022cogvideo , where image tokenizers are directly applied without temporal modeling.
-Block-based tokenization yan2021videogpt ; ge2022tats ; yu2023magvit improves on this by using 3D convolutions to encode short clips with local spatio-temporal context.
-More recently, 1-D tokenizers wang2024larp introduce learnable holistic tokens to capture global information across the entire video.
-Despite these advances, these methods are inherently non-causal, relying on full access to the video during encoding, which limits their applicability to streaming or autoregressive generation tasks.
+To handle continuous visual signals, VQ-VAE (Van Den Oord et al., 2017) and its extensions (Esser et al., 2021; Yu et al., 2022; Zheng & Vedaldi, 2023; Zhang et al., 2023; Mentzer et al., 2023) introduce vector quantization to convert images into discrete 2D code sequences.
+Following works such as TiTok (Yu et al., 2024b) further propose a 1D tokenizer that tokenizes images into compact 1D latent sequences.
+When applied to videos, early approaches use per-frame tokenization (Gupta et al., 2022; Wu et al., 2022; Hong et al., 2023), where image tokenizers are directly applied without temporal modeling.
+Block-based tokenization (Yan et al., 2021; Ge et al., 2022; Yu et al., 2023) improves on this by using 3D convolutions to encode short clips with local spatio-temporal context.
+More recently, LARP (Wang et al., 2024a) introduces holistic tokens to capture global video information.
+However, similar to most prior approaches, its representations remain non-causal and fixed in length, which limits their applicability to streaming or autoregressive generation tasks.
+In contrast, we target causal, adaptive video tokenizer that represents videos with variable‑length latent tokens, enabling online, low-latency processing while providing flexible representation granularity for downstream tasks.
 To address the limitations of non-causal tokenizers, several works have explored causal temporal tokenization architectures.
-One line of work adopts causal convolutions for autoregressive modeling yu2023magvitv2 .
-Building on this, Cosmos-Tokenizer agarwal2025cosmos introduces both causal convolution layers and causal temporal attention layers for temporal modeling.
-In parallel, another line of work focuses on transformer-based architectures villegas2022phenaki ; wang2024omnitokenizer , where causality is enforced via attention masks, enabling autoregressive modeling using only past frames.
-Following this trend, our method also adopts a transformer-based causal architecture, facilitating better adaptation to downstream Physical AI applications that require a temporal causal setting.
+One line of work adopts causal convolutions for autoregressive modeling (Yu et al., 2024a).
+Building on this, Cosmos-Tokenizer (NVIDIA, 2025) introduces causal convolution layers and causal temporal attention layers for temporal modeling.
+In parallel, another line of work focuses on transformer-based architectures (Villegas et al., 2022; Wang et al., 2024b), where causality is enforced via attention masks, enabling autoregressive modeling using only past frames.
+Following such trend, we also adopt a transformer-based causal architecture.
 2.2 Adaptive Tokenizer
-Recent works have explored adaptive tokenization to allow variable length representations.
-For image tokenization, ALIT duggal2024alit dynamically adjusts the token count by recursively distilling 2D image tokens into 1D latent sequence.
-Another common strategy is “tail drop” koike2020stochastic , which applies higher dropout rate on the tail of the latent feature to learn ordered representations.
-Building on this idea, One-D-Piece miwa2025onedpiece and FlexTok bachmann2025flextok introduce tail-drop regularization for 1D image tokenizers.
-To support adaptive inference, CAT shen2025cat predicts image complexity using captions and large language models (LLMs), assigning each input to one of three compression ratios (8$\times$, 16$\times$, 32$\times$).
-Extending adaptive tokenization to videos, ElasticTok yan2024elastictok applies tail dropping at the end of each block and supports adaptive inference by selecting token counts via a fixed threshold.
+[…]
+Extending adaptive tokenization to videos, ElasticTok (Yan et al., 2024) applies tail dropping at the end of each block and supports adaptive inference by selecting token counts via a fixed threshold.
 However, due to the local spatial-temporal nature of these tokens—rather than forming a unified 1D sequence—it is difficult to gather informative content at the head, limiting the effectiveness of tail dropping.
@@
 3.1.2 Block Causal Transformer
+The block causal transformer employs a causal encoder to extract block-wise representations, a block-mask sampler to randomly sample tokens, an SVQ (Wang et al., 2024a) quantizer for discretization, and a causal decoder to reconstruct videos.
 Block Causal Encoder.
-To design an adaptive tokenizer, it is crucial to avoid tokens that explicitly correspond to local spatial-temporal regions.
-To this end, the encoder employs a set of learnable latent tokens to transform the patch embeddings into a 1-D sequence.
-Specifically, the latent tokens $q_{\mathrm{enc}}\in\mathbb{R}^{N\times d}$ are concatenated with image tokens $e\in\mathbb{R}^{L\times d}$, forming a combined input of $(L+N)$ tokens to a block-causal transformer encoder $\mathcal{E}$.
-The output corresponding to the latent tokens, denoted as $\tilde{z}=\mathcal{E}(e\oplus q_{\mathrm{enc}})_{L:(L+N)}$, serves as the learned latent representation.
-Across blocks, block-causal attention is applied: all tokens in a block (both latent and image tokens) can only attend to tokens from the same or previous blocks, thereby enforcing a causal dependency structure.
+To design an adaptive tokenizer, it is crucial to disentangle the positional dependencies between the latent space and the local spatiotemporal regions.
+Motivated by Yu et al.
+(2024b); Wang et al.
+(2024a), the encoder employs learnable latent tokens to transform patch embeddings into a 1D sequence.
+Specifically, the latent tokens $q_{\mathrm{enc}}\in\mathbb{R}^{N\times d}$ are concatenated with video tokens $e\in\mathbb{R}^{L\times d}$, forming a combined input of $(L+N)$ tokens to a block-causal transformer encoder $\mathcal{E}$.
+The output corresponding to the latent tokens, denoted as $\tilde{z}=\mathcal{E}(e\oplus q_{\mathrm{enc}})_{L:(L+N)}$, serves as the latent representation.
+Across blocks, block-causal attention is applied: all tokens in a block (both latent tokens $q_{\mathrm{enc}}$ and video tokens $e$) can only attend to tokens from the same or previous blocks, thereby enforcing a causal dependency structure.
 The pattern for the encoder attention mask is shown in Fig.
@@
 Block-mask Sampler.
-We train AdapTok to learn variable length codes over a short video block (e.g., 4 frames) by randomly dropping the tail codes for each block during training.
+Inspired by  Miwa et al.
+(2025); Bachmann et al.
+(2025); Yan et al.
+(2024), we train AdapTok to learn variable-length codes over a short video block (e.g., 4 frames) by randomly dropping the tail codes for each block during training.
 Specifically, for each latent block $\tilde{z}_{i}$, we randomly sample the number of tokens to retain, denoted as $\ell_{i}$.
@@
 The mask $m^{\prime}$ is then applied to the latent token sequence $\tilde{z}$, where tokens with $m_{j}^{\prime}=0$ are dropped during training.
-The resulting sequence, denoted as $z=\mathrm{BlockTailDrop}(\tilde{z}\ |\ m^{\prime})$, is used for decoding, enabling the model to learn efficient, variable-length representations.
-Quantizer.
-Following wang2024larp , a stochastic vector quantization (SVQ) is adopted to the quantizer $\mathcal{Q}$.
-Similar to VQ, a codebook $C\in\mathbb{R}^{c\times d^{\prime}}$ containing $c$ codes is maintained.
-Given a visual feature $z$, the cosine similarity with all code vectors in $C$ is computed, followed by a softmax to obtain a probability distribution.
-An index $x_{\mathrm{Ind}}$ is then sampled from this distribution via a categorical distribution:
-$x_{\mathrm{Ind}}\sim\mathrm{Categorical}\left(\mathrm{softmax}\left(\left\{\frac{v\cdot C_{i}}{\|v\|\|C_{i}\|}\right\}_{i=1}^{c}\right)\right).$ (2)
-Given the sampled index $x_{\mathrm{Ind}}$, the quantized feature $z_{q}$ is retrieved via codebook lookup, i.e., $z_{q}=C_{x_{\mathrm{Ind}}}$.
-To enable differentiable training, the straight-through estimator bengio2013estimating is applied.
+The resulting sequence, denoted as $z=\mathrm{BlockTailDrop}(\tilde{z}\ |\ m^{\prime})$, is then discretized into the quantized feature $z_{q}$.
 Block Causal Decoder.
-The quantized tokens are then concatenated with a set of learnable image latent tokens $q_{\mathrm{dec}}\in\mathbb{R}^{L\times d}$, and the combined $(N^{\prime}+L)$ tokens are fed into a block-causal decoder $\mathcal{D}$, implemented as a transformer.
-The last $L$ outputs, denoted as $\hat{z}=\mathcal{D}(q_{\mathrm{dec}}\oplus z_{q},\mathcal{M}^{\prime})_{N^{\prime}:(N^{\prime}+L)}$, are used to reconstruct the video $\hat{x}\in\mathbb{R}^{T\times H\times W\times 3}$.
-Here, $N^{\prime}$ denotes the number of latent tokens remaining after tail-drop, and $\mathcal{M}^{\prime}$ is the attention mask applied in the decoder.This mask is derived by applying the dropout mask $m^{\prime}$ to the original block-causal attention pattern $\mathcal{M}$, as shown in Fig 2(b-ii),
+The quantized tokens $z_{q}$ are concatenated with a set of learnable video latent tokens $q_{\mathrm{dec}}\in\mathbb{R}^{L\times d}$, and the combined $(N^{\prime}+L)$ tokens are fed into a block-causal decoder $\mathcal{D}$, implemented as a transformer.
+The last $L$ outputs, denoted as $\hat{z}=\mathcal{D}(q_{\mathrm{dec}}\oplus z_{q},\mathcal{M}^{\prime})_{N^{\prime}:(N^{\prime}+L)}$, are used to reconstruct the video.
+Here, $N^{\prime}$ denotes the number of quantized tokens remaining after tail-drop, and $\mathcal{M}^{\prime}$ is the attention mask applied in the decoder.This mask is derived by applying the dropout mask $m^{\prime}$ to the original block-causal attention pattern $\mathcal{M}$, as shown in Fig 2(b-ii),
 $\mathcal{M}^{\prime}=\mathrm{BlockTailDrop}(\mathcal{M}\ |\ m^{\prime}).$ (2)
-Within this mask, quantized latent tokens $z_{q}$ attend only to quantized tokens from the same or preceding blocks, while image latent tokens $q_{\mathrm{dec}}$ attend to all image tokens within the same block and to quantized tokens from the same or earlier blocks.
+Within this mask, quantized tokens $z_{q}$ attend only to quantized tokens from the same or preceding blocks, while video latent tokens $q_{\mathrm{dec}}$ attend to all video tokens within the same block and to quantized tokens from the same or earlier blocks.
 3.1.3 Training
-Following previous works esser2021vqgan ; ge2022tats ; wang2024larp , the tokenizer is optimized with a composite objective:
+Following previous works Esser et al. (2021); Ge et al. (2022); Wang et al. (2024a), the tokenizer is optimized with a composite objective:
 $\mathcal{L}=\mathcal{L}_{R}+\mathcal{L}_{VQ}+\mathcal{L}_{P}+\mathcal{L}_{G}+\mathcal{L}_{prior},$ (3)
-where $\mathcal{L}_{R}$ denotes the $L_{1}$ reconstruction loss, $\mathcal{L}_{VQ}$ the quantizer loss, $\mathcal{L}_{P}$ the perceptual loss zhang2018unreasonable , $\mathcal{L}_{G}$ the adversarial loss goodfellow2014generative to enhance visual quality, and $\mathcal{L}_{prior}$ the autoregressive prior loss wang2024larp for modeling the latent sequence.
+where $\mathcal{L}_{R}$ denotes the $L_{1}$ reconstruction loss, $\mathcal{L}_{VQ}$ the quantizer loss, $\mathcal{L}_{P}$ the perceptual loss (Zhang et al., 2018), $\mathcal{L}_{G}$ the adversarial loss (Goodfellow et al., 2014) to enhance visual quality, and $\mathcal{L}_{prior}$ the autoregressive prior loss (Wang et al., 2024a) for modeling the latent sequence.
 3.2 Adaptive Scorer
@@
 At each iteration, a target block index $q\sim\mathcal{U}(0,K-1)$ is sampled.
-We duplicate the latent sequence ${z}$ and apply a series of latent masks $\{m^{\prime}_{p}\}$, each corresponding to a different number of tokens $\ell_{q}$, ranging from the supported range of encoding lengths.
+We duplicate the latent sequence z{z} and apply a series of latent masks $\{m^{\prime}_{p}\}$, each corresponding to a different number of tokens $\ell_{q}$, ranging from the supported range of encoding lengths.
 The token lengths $\ell_{i<q}$ for preceding blocks are randomly sampled, while all subsequent blocks are fully masked, i.e., $\ell_{i>q}=0$.
@@
 $m_{p}^{\prime}=\ [m_{p,1}\oplus m_{p,2}\oplus\cdots\oplus m_{p,q}\oplus\mathbf{0}\ ],\ m_{p,i<q}=\ [\mathbbm{1}_{j\leq\ell_{i}}]_{j=1}^{M},\ m_{p,q}=\ [\mathbbm{1}_{j\leq p}]_{j=1}^{M}.$ (4)
-For each $m^{\prime}_{p}$, the perceptual loss $\mathcal{L}_{P}$ over the $q$-th block is computed, serving as the quality scores $s$.
+For each $m^{\prime}_{p}$, the perceptual loss $\mathcal{L}_{P}$ over the $q$-th block is computed, used as the quality scores $s$.
 Block Causal Scorer.
@@
 The dropout mask $m^{\prime}_{s}$ is constructed in a manner similar to $m^{\prime}_{p}$ (used for ground-truth generation), except that $\ell_{q}$ is fixed to the maximum block token length $M$.
-The mean squared error (MSE) loss $\mathcal{L}_{s}$ is computed between the predicted score $\hat{s}$ and the ground-truth quality scores $s$.
+The MSE loss $\mathcal{L}_{s}$ is computed between the predicted score $\hat{s}$ and the ground-truth quality scores $s$.
 3.3 Adaptive Token allocation
@@
 The optimal assignment $b^{*}$ yields the final token count for each sample: $n_{k}=\sum_{j}j\cdot b^{*}_{kj}$.
-1 Inputs: Video $x$, Block idx $q$;
+1Inputs: Video $x$, Block idx $q$;
 2 Hyperparameters: Average Token $N_{b}$, tokens per block $M$, batch size $B$;
@@
 Given a sequence of adaptively allocated tokens from the scorer with IPAL, an end-of-block token <EOB> is appended to the end of each block.
-The resulting concatenated sequence is then fed into a Llama-style transformer touvron2023llama ; touvron2023llama2openfoundation for autoregressive generation.
+The resulting concatenated sequence is then fed into a Llama-style transformer (Touvron et al., 2023a; b) for autoregressive generation.
 Formally, let $y=(y_{1},y_{2},\cdots,y_{S})$ denote the multi-block adaptive sequence.
@@
 where $c$ denotes the conditions, e.g., class labels for class-conditional generation or context tokens for frame prediction, and $\theta$ represents the trainable parameters of the transformer.
+Table 1: Comparison of Video Reconstruction FVD on UCF-101.
+All models use a causal visual tokenizer.
+Bold and underline indicate the best and the second best performance, respectively.
+$\dagger$ denotes models trained using the same dataset and receipe as ours.
+Method Data size Codebook Tokens rFVD $\downarrow$
+OmniTokenizer (Wang et al., 2024b) 1.4M 8,192 1,280 42
+ElasticTok (Yan et al., 2024) 356M 64,000 1,024 390
+ElasticTok (Yan et al., 2024) 356M 64,000 2,048 93
+Cosmos-Tokenizer-DV (NVIDIA, 2025) 100M 64,000 1,280 140
+OmniTokenizer $\dagger$ (Wang et al., 2024b) <<0.5M 8,192 1,280 94
+ElasticTok $\dagger$ (Yan et al., 2024) <<0.5M 64,000 1,022 230
+CausalTok $\dagger$ <<0.5M 8,192 1,024 37
+AdapTok (Ours) <<0.5M 8,192 512 60
+AdapTok (Ours) <<0.5M 8,192 1,024 36
+AdapTok (Ours) <<0.5M 8,192 2,048 28
+Table 2: Class-conditional generation results on UCF-101 and frame prediciton results on Kinetics-600. $\ddagger$ denotes using pretrained image tokenizer.
+gFVD $\downarrow$
+Method Data size Params K600 UCF
+Phenaki (Villegas et al., 2022) 465M 1.8B 36.4 /
+MAGVIT-AR (Yu et al., 2023) 12M 306M / 265
+MAGVIT-v2-AR (Yu et al., 2024a) 1.8M 840M / 109
+OmniTokenizer (Wang et al., 2024b) 1.4M 650M 32.9 191
+CogVideo $\ddagger$ (Hong et al., 2023) 5.4M 9.4B 109.2 626
+Video-LaVIT $\ddagger$ (Jin et al., 2024) 10M 7B / 281
+TATS (Ge et al., 2022) <<0.5M 321M / 332
+CausalTok <<0.5M 633M / 80
+AdapTok-AR (Ours) <<0.5M 633M 11 67
 4 Experiments
@@
 Datasets.
-We train the model and evaluate its performance on two standard datasets: video reconstruction and class-conditional generation on UCF-101 soomro2012ucf101 and frame prediction on Kinetics-600 carreira2018k600 with 5-frame conditioning.
+We train the model and evaluate its performance on two standard datasets: video reconstruction and class-conditional generation on UCF-101 (Soomro et al., 2012) and frame prediction on Kinetics-600 (Carreira et al., 2018) with 5-frame conditioning.
 Implementation details.
-Following previous works ge2022tats ; yu2023magvitv2 , we adopt $16\times 128\times 128$ video clips for both training and evaluation across all tasks.
+Following Ge et al.
+(2022); Yu et al.
+(2024a), we adopt $16\times 128\times 128$ video clips for both training and evaluation.
 Patch embeddings are extracted with a patch size of $4\times 8\times 8$, resulting in $L=1024$ tokens per video.
 A sequence of latent tokens $q_{\mathrm{enc}}$ is used to extract these patch embeddings, and is partitioned into $K=4$ blocks, each containing $M=512$ tokens.
-These latent tokens are then quantized using a codebook $C\in\mathbb{R}^{8192\times 16}$.
 For adaptive token sampling, we employ a block-wise mask sampler.
-In each block, the number of tokens is sampled from a truncated Gaussian distribution with mean $\mu=256$, standard deviation $\sigma=128$, and bounded between $M_{\mathrm{min}}=32$ and $M_{\mathrm{max}}=512$.
+In each block, the token number is sampled from a truncated Gaussian distribution with mean $\mu=256$, standard deviation $\sigma=128$, and bounded between $M_{\mathrm{min}}=32$ and $M_{\mathrm{max}}=512$.
 AdapTok is trained on UCF-101 and Kinetics-600 for 250 epochs with a batch-size of 128.
-Adam kingma2014adam is adopt as the optimizer with hyperparameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$.
-The learning rate is linearly warmed up to $10^{-4}$ and subsequently decayed to $10^{-6}$ using a cosine scheduler.
-Fréchet Video Distance (FVD) thomas2018fvd is adopt as the main evaluation metric for both reconstruction and generation. Additionally, we report MSE, PSNR and LPIPS for video reconstruction.
+Adam (Kingma & Ba, 2014) is adopt as the optimizer with hyperparameters $\beta_{1}=0.5$ and $\beta_{2}=0.9$.
+The learning rate is linearly warmed up to $10^{-4}$ and then decayed to $10^{-6}$ using a cosine scheduler.
+Fréchet Video Distance (FVD) (Unterthiner et al., 2018) is adopt as the primary metric for reconstruction and generation. Additionally, we report PSNR and LPIPS for video reconstruction.
 4.2 Main Results
@@
 We first evaluate the video reconstruction performance of AdapTok on UCF-101, with results summarized in Tab.
-2.
-We can observe that, even with fewer training data and latent tokens, AdapTok significantly outperforms existing causal video tokenizers, achieving an rFVD of 28 with 2048 latent tokens and 36 with 1024 tokens.
-To further evaluate efficiency under varying token budgets, we compare AdapTok with several causal tokenizer baselines.
+1.
 In particular, we implement a strong baseline named CausalTok that shares the same block-causal architecture as AdapTok, but applies a standard non-adaptive pipeline.
-As shown in Fig.
-3, AdapTok consistently achieves lower rFVD across different token counts.
-Remarkably, it achieves an FVD of 60 using only 512 tokens, outperforming most baselines, and achieves comparable performance to CausalTok while using approximately 1.8$\times$ fewer tokens.
+To ensure fair comparison, we also reproduced previous works (i.e., ElasticTok and OmniTokenizer) using the same dataset and receipe as ours.
+With fewer training data and latent tokens, AdapTok significantly outperforms existing causal video tokenizers, achieving an rFVD of 28 with 2048 tokens and 36 with 1024 tokens.
+[…]
 Video Generation.
@@
 2, our model achieves competitive performance on both class-conditional generation and frame prediction.
-Notably, with only 633M parameters, AdapTok-AR achieves a gFVD of 11 on Kinetics-600 and 67 on UCF-101.
-Table 1: Comparison of Video Reconstruction FVD on UCF-101. All models use a causal visual tokenizer. Bold and underline indicate the best and the second best performance, respectively.
-Method Data size Codebook Tokens rFVD $\downarrow$
-OmniTokenizer wang2024omnitokenizer 1.4M 8,192 1,280 42
-ElasticTok yan2024elastictok 356M 64,000 1,024 390
-ElasticTok yan2024elastictok 356M 64,000 2,048 93
-Cosmos-Tokenizer-DV agarwal2025cosmos 100M 64,000 1,280 140
-AdapTok (Ours) <0.5M 8,192 512 60
-AdapTok (Ours) <0.5M 8,192 1,024 36
-AdapTok (Ours) <0.5M 8,192 2,048 28
-Table 2: Class-conditional generation results on UCF-101 and frame prediciton results on Kinetics-600. $\ddagger$ denotes using pretrained image tokenizer.
-gFVD $\downarrow$
-Method Data size Params K600 UCF
-Phenaki villegas2022phenaki 465M 1.8B 36.4 /
-MAGVIT-AR yu2023magvit 12M 306M / 265
-MAGVIT-v2-AR yu2023magvitv2 1.8M 840M / 109
-OmniTokenizer wang2024omnitokenizer 1.4M 650M 32.9 191
-CogVideo hong2022cogvideo  $\ddagger$ 5.4M 9.4B 109.2 626
-Video-LaVIT jin2024videolavit  $\ddagger$ 10M 7B / 281
-TATS ge2022tats <0.5M 321M / 332
-AdapTok-AR (Ours) <0.5M 633M 11 67
+With only 633M parameters, AdapTok-AR achieves a gFVD of 11 on Kinetics-600 and 67 on UCF-101.
+Notably, with similar reconstruction performance, AdapTok achieves better generation results than CausalTok, demonstrating the advantage of adaptively allocating tokens for different videos.
+Model Scaling.
+To further investigate the impact of model size on performance, we trained models across 3 different sizes: AdapTok-S, AdapTok-L, and AdapTok-XL.
+As shown in Tab.
+3, the reconstruction quality improves consistently with increasing model size.
+Such results indicate that AdapTok benefits from the increased model capacity, highlighting the scaling potential of our method.
+Table 3: Performance of AdapTok with different model sizes. Continuous improvement is achieved as the model size increases.
+Model Params rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
+AdapTok-S 59M 87.32 24.23 0.151
+AdapTok-L 259M 36.36 25.72 0.144
+AdapTok-XL 913M 32.43 26.29 0.103
 4.3 Ablation Studies
@@
 As shown in Tab.
-3a, leveraging both mechanisms results in the best performance, achieving the lowest rFVD of 36.36.
+4, leveraging both mechanisms results in the best performance, achieving the lowest rFVD of 36.36.
 The performance gain becomes more pronounced with lower token budgets in inference, highlighting the importance of our adaptive mechanisms for efficient video representation.
-Table 3: Ablation studies. (a) Ablation on adaptive training and inference paradigms. (b) Evaluation on different token allocation strategies. (c) Comparison on token allocation strategies.
-Tokens Mask Sampler Scorer rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
+Table 4: Ablation on adaptive training and inference.
+Tokens Sampler Scorer rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
 1024 37.13 25.92 0.111
@@
 59.96 24.06 0.144
-(a) Adaptive paradigms.
+[…]
+Method Time (ms/video)
+ElasticTok 571.7
+AdapTok 50.9
+Table 6: Ablation on token allocation strategies.
 Method rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
 ILP 36.36 25.72 0.114
-(b) Token allocation strategies.
+Table 7: Comparison on scoring metrics.
 Metric rFVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$
@@
 Perceptual 36.28 25.72 0.113
-(c) Scoring metrics.
 Token allocation strategies.
@@
 As reported in Tab.
-3b, the ILP strategy outperforms the other methods across all metrics.
-For a comparison under varying token lengths, refer to Fig.
-4.
-Figure 4: Comparison of token allocation strategies. ILP achieves the best performance across four metrics: (a) FVD, (b) PSNR, (c) MSE, and (d) LPIPS.
+6, the ILP strategy outperforms the other methods across all metrics.
+Fig.
+4 further compares the different token allocation strategies across varying token lengths, demonstrating that the ILP-based allocation strategy consistently outperforms the alternatives.
+Figure 4: Comparison of token allocation strategies. ILP achieves the best performance across all metrics: (a) FVD, (b) PSNR, and (c) LPIPS.
 Token allocation scoring metrics.
@@
 As shown in Tab.
-3c and Fig.
-5, each scoring metric performs best on its corresponding evaluation metric, e.g., PSNR score yields the highest PSNR (Fig.
-5b), MSE score leads to the lowest MSE (Fig.
-5c).
-Notably, using perceptual loss as the scoring metric also leads to superior performance on other evaluation metrics, such as rFVD (Fig.
-5a), indicating its stronger correlation with overall perceptual quality.
-Figure 5: Comparison of scoring metrics. Perceptual loss achieves better overall performance, especially on FVD and LPIPS.
+7, each scoring metric performs best on its corresponding evaluation metric, e.g., PSNR score yields the highest PSNR.
+Notably, using perceptual loss as the scoring metric also leads to superior performance on other evaluation metrics, such as rFVD, indicating its stronger correlation with overall perceptual quality.
+For a comparison under varying token lengths, please refer to the Appendix C.
+[…]
+Figure 5: Attention maps for latent tokens. Each map shows the attention distribution of a token over spatial patches, with tokens ordered from top-left to bottom-right.
 4.4 Visualizations
-1D latent token space matters.
-In Fig.
-6, we visualize reconstruction results under different token budgets.
-Compared to ElasticTok, which relies on local 2D spatial tokens and tends to produce block-like artifacts, AdapTok leverages 1D latent space that enables a coarse-to-fine reconstruction process: early tokens capture global structure, while later tokens refine local details.
-Figure 6: Reconstruction comparison with ElasticTok yan2024elastictok under varying token budgets.
-AdapTok reconstructs videos in a coarse-to-fine manner, where early tokens capture global structure and later ones refine local details, while ElasticTok suffers from blocky artifacts.
 Head tokens encode global information.
 The tail token drop strategy encourages the head tokens to capture more global information.
-As shown in Fig.7, attention maps reveal that early tokens attend broadly across the entire spatial field, indicating their role in modeling global semantics.
-In contrast, mid-to-late tokens focus on localized regions, refining finer details.
-Figure 7: Attention maps for latent tokens.
-Each map shows the attention distribution of a token over spatial patches, with tokens ordered from top-left to bottom-right.
-Early tokens capture global context, while later tokens focus on local details.
+As shown in Fig.5, early tokens capture global context, while later tokens focus on local details.
+Please refer to the Appendix D for more visualization results.
 5 Conclusion
@@
 In this work, we focus on the design of a sample-wise, content-aware, and temporally adaptive tokenization framework.
-Our current implementation adopts a discrete tokenizer based on VQ-VAEs van2017vqvae ; esser2021vqgan , and future work may explore continuous alternatives to validate the generality of the proposed framework.
+Our current implementation adopts a discrete tokenizer based on VQ-VAEs (Van Den Oord et al., 2017; Esser et al., 2021), and future work may explore continuous alternatives to validate the generality of the proposed framework.
 In addition, our model is trained on less than 0.5M publicly available videos due to limited computational resources.
 Future efforts will focus on scaling up model capacity and dataset diversity to improve generalization and applicability across broader domains.
+Ethics Statement
+This work uses only publicly available datasets that have been rigorously filtered to mitigate potential biases and ethical concerns.
+Reproducibility Statement
+The implementation details of our method are fully described in Sec.
+4.1 and Appendix.
+A, including both hyper-parameter settings and training costs.
+Additionally, the pseudo-code for IPAL and details of other token allocation strategies are provided in Algorithm 1 and Appendix.
+B.
+The source code and checkpoints will be released to enable reproduction of the main results presented in this paper.
 Appendix A Implementation Details
-The detailed training hyper-parameter settings for the AdapTok, Scorer, AdapTok-AR (class-conditional generation on UCF-101) and AdapTok-FP (frame prediciton on Kinetics-600) are reported in Tab. 4.
+The detailed training hyper-parameter settings for the AdapTok, Scorer, AdapTok-AR (class-conditional generation on UCF-101) and AdapTok-FP (frame prediciton on Kinetics-600) are reported in Table 8.
+The architectural configurations of the different AdapTok variants are listed in Table 9.
 Table 8: Hyper-parameters for AdapTok models.
@@
 Model parameters
-[…]
+Parameters 195M 89M 633M 633M
 Frame Resolution $16\times 128\times 128$ $16\times 128\times 128$ $16\times 128\times 128$ $16\times 128\times 128$
@@
 Training
-[…]
+Optimizer Adam Adam AdamW AdamW
 Learning rate $1e^{-4}$ $1e^{-4}$ $6e^{-4}$ $6e^{-4}$
@@
 Training Time 112h 9h 59h 55h
-Appendix B Adaptive Inference
+Table 9: Model configurations of AdapTok variants.
+Model Hidden Size Depth Heads Parameters
+AdapTok-S 512 6 8 59M
+AdapTok-L 768 12 12 259M
+AdapTok-XL 1024 24 16 913M
+Appendix B More Technical Details
+B.1 Quantization method
+Following Wang et al.
+(2024a), a stochastic vector quantization (SVQ) is adopted to the quantizer $\mathcal{Q}$.
+Similar to VQ, a codebook $C\in\mathbb{R}^{c\times d^{\prime}}$ containing $c$ codes is maintained.
+Given a visual feature $z$, the cosine similarity with all code vectors in $C$ is computed, followed by a softmax to obtain a probability distribution.
+An index $x_{\mathrm{Ind}}$ is then sampled from this distribution via a categorical distribution:
+$x_{\mathrm{Ind}}\sim\mathrm{Categorical}\left(\mathrm{softmax}\left(\left\{\frac{v\cdot C_{i}}{\|v\|\|C_{i}\|}\right\}_{i=1}^{c}\right)\right).$ (8)
+Given the sampled index, the quantized feature $z_{q}$ is retrieved from the codebook, i.e., $z_{q}=C_{x_{\mathrm{Ind}}}$.
+To enable differentiable training, the straight-through estimator Bengio et al.
+(2013) is applied.
+B.2 Adaptive Inference
 In addition to the Integer Linear Programming (ILP) method detailed in the main paper, we provide further details for the other three token allocation methods, including Fixed, BiThr and BiDelta.
@@
 The search aims to find a threshold over these deltas such that the resulting token allocation meets the desired token count.
-1 Inputs: Video $x$, Block idx $i$;
+1Inputs: Video $x$, Block idx $i$;
 2 Hyperparameters: Average Token $N_{b}$, tokens per block $M$, max iterations $K$;
@@
 Algorithm 2  Binary Search
-Appendix C Visualizations
-C.1 Video Reconstruction
+Appendix C Additional Experiments
+Ablation on token allocation scoring metrics.
+Fig.
+6 presents detailed comparisons of token allocation under varying token lengths for different scoring metrics.
+Each metric achieves the best performance on its corresponding evaluation metric, while perceptual loss consistently yields strong results across multiple metrics, demonstrating its effectiveness for guiding adaptive token allocation.
+Figure 6: Comparison of scoring metrics. Using perceptual loss as the scoring metric achieves better overall performance, especially on FVD and LPIPS.
+Further analysis on computational costs.
+Although IPAL is based on ILP, its actual runtime overhead is not significant.
+Table 10-12 shows the runtime scalability of IPAL w.r.t.
+mini-batch size, the number of blocks and average token numbers.
+Specifically, across batch sizes ranging from 8 to 1024, FVD remains stable between 36 and 37, while IPAL accounts for only about 15% of total inference time.
+The allocation time also remains stable with varying average token numbers, while it increases moderately as the number of blocks grows.
+These results suggest that our method is efficient, scalable, and practical for real-world deployment scenarios.
+Table 10: Ablation on mini-batch size with an average token of 512.
+Batch size FVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Time (ms/video) IPAL Time
+Total IPAL Proportion (%)
+8 60.39 24.01 0.146 66.3 28.0 42.2
+16 60.14 24.04 0.145 53.7 16.7 31.2
+64 59.96 24.06 0.144 45.2 8.6 19.1
+128 60.61 24.07 0.144 44.1 7.0 15.8
+256 61.48 24.08 0.144 44.0 6.7 15.2
+512 61.44 24.08 0.144 43.0 6.3 14.8
+1024 61.37 24.09 0.144 44.1 6.8 15.5
+Table 11: Ablation on mini-batch size with an average token of 1024.
+Batch size FVD $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ Time (ms/video) IPAL Time
+Total IPAL Proportion (%)
+8 36.55 25.68 0.115 69.0 28.1 40.7
+16 36.80 25.71 0.114 56.8 16.9 29.8
+64 36.36 25.72 0.114 48.1 8.5 17.8
+128 36.52 25.73 0.114 46.7 7.2 15.4
+256 37.02 25.74 0.114 47.4 6.9 14.6
+512 36.93 25.74 0.113 46.9 7.0 14.8
+1024 37.01 25.75 0.113 47.3 7.5 15.8
+Table 12: Runtime ablation of IPAL on block and token counts. (a) Ablation on block numbers. (b) Ablation on average token counts.
+Blocks Tokens IPAL Time (ms/video)
+1 1024 7.9
+2 1024 7.1
+4 1024 8.0
+8 1024 10.6
+(a) Number of blocks.
+Blocks Tokens IPAL Time (ms/video)
+4 32 7.8
+4 128 8.5
+4 512 8.3
+4 1024 8.0
+(b) Average tokens counts.
+Appendix D More Visualizations
+D.1 Video Reconstruction
+1D latent token space matters.
+In Fig.
+7, we visualize reconstruction results under different token budgets.
+Compared to ElasticTok, which relies on local 2D spatial tokens and tends to produce block-like artifacts, AdapTok leverages 1D latent space that enables a coarse-to-fine reconstruction process: early tokens capture global structure, while later tokens refine local details.
+Figure 7: Reconstruction comparison with ElasticTok Yan et al.
+(2024) under varying token budgets.
+AdapTok reconstructs videos in a coarse-to-fine manner, where early tokens capture global structure and later ones refine local details, while ElasticTok suffers from blocky artifacts.
 Content-aware allocation.
@@
 Figure 9: AdapTok adaptively allocates tokens based on temporal dynamics. Red boxes highlight the scene transitions, where AdapTok allocates more tokens to capture important temporal changes.
-C.2 Video Generation
+D.2 Video Generation
 We present class conditional generation results on UCF-101 in Fig. 10 and frame prediction results on Kinetics-600 in Fig. 11.
@@
 Figure 11: Frame prediction results on Kinetics-600. Red boxes represent the condition frames.
-C.3 Token Contribution Analysis
+D.3 Token Contribution Analysis
 Fig.
@@
 Early tokens attend broadly to capture global context, while later ones focus on local details.
-C.4 Scorer Predictions
+D.4 Scorer Predictions
 Fig.
@@
 (b) Corresponding reconstruction results of frames from the first block under different token counts.
+Appendix E The Use of Large Language Models (LLMs)
+We used LLMs solely to assist with grammar correction of the submission. All research ideas, experiments, and analyses presented in this work were conceived, conducted, and verified by the authors.